OpenStudy (fanduekisses):

[DID I SOLVE THIS CORRECTLY????] Related Rates: a 20 foot ladder slides down a wall at 5. at what speed is the bottom slidng out when the top is 1o feet from the floor

OpenStudy (fanduekisses):

I keep getting \[5\sqrt{3}\]

OpenStudy (fanduekisses):

but it's none of my answer choices :'(

OpenStudy (fanduekisses):

\[2a \frac{ da }{ dt }+2b \frac{ db }{ dt }= 0\] \[2(10)(-5)+2(10\sqrt(3))\frac{ db }{ dt }\]

OpenStudy (freckles):

I like to draw a picture first.